OpenStudy (anonymous):

5^(x-2)25^(4x)=(1)/(√(125^(2x+3))) Find x

OpenStudy (anonymous):

x=0.069

OpenStudy (anonymous):

how to do?

OpenStudy (anonymous):

\[(x-2)*25^{4x}\log5_{5} = (-1/2) * (2*x+3)*\log125_{5}; (x-2)*25^{4x}=-3*x-4.5\]

OpenStudy (anonymous):

erm i dont know how to do logarithm yet.

OpenStudy (anonymous):

without him nothing